You are the owner of a local Honda dealership. Unlike other dealerships in the area, you take pride in your “No Haggle” sales policy. Last year, your dealership earned record profits of $1.2 million. In your market, you compete against two other dealers, and the market-level price elasticity of demand for midsized Honda automobiles is -1.4. In each of the last five years, your dealership has sold more midsized automobiles than any other Honda dealership in the nation. This entitled your dealership to an additional 35 percent off the manufacturer’s suggested retail price (MSRP) in each year. Taking this into account, your marginal cost of a midsized automobile is $11,000.

What price should you charge for a midsized automobile if you expect to maintain your record sales?
Answer #1

Lerner Index = -1 / Elasticity of demand = (Price - MC) / Price

- 1 / -1.4 = (P - 11,000) / P

1 / 1.4 = (P - 11,000) / P

P = 1.4P - 15,400

0.4P = 15,400

P = $38,500